OpenStudy (anonymous):

interesting integral:

OpenStudy (anonymous):

part 1 let\[I(a,b) = \int\limits^1_0x^a(1-x)^bdx\] show that \[I(a,b) = I(b,a)\]

OpenStudy (blockcolder):

That reminds me of an exercise in Stewart's.

OpenStudy (anonymous):

part 2 show that \[I(a,b) = I(a+1,b) + I(a,b+1)\]

OpenStudy (anonymous):

part 3 show that \[(a+1)I(a, b) = bI(a+1, b−1)\] when a and b are positive. hence evaluate I(a,b) when a and b are positive integers

OpenStudy (blockcolder):

For part I, let u=1-x: \[\int_0^1x^a(1-x)^b\ dx=-\int_1^0(1-u)^au^b\ du=\int_0^1 (1-x)^ax^b\ dx\]

OpenStudy (anonymous):

yeeeah

OpenStudy (blockcolder):

I'm a little stuck on part 2. Does it require integration by parts?

OpenStudy (blockcolder):

My bad, it doesn't require integration by parts. \[\int_0^1 x^{a+1}(1-x)^b\ dx+\int_0^1 x^a(1-x)^{b+1}\ dx\\ =\int_0^1 x^{a+1}(1-x)^b+x^a(1-x)^{b+1}\ dx\\ =\int_0^1 x^a(1-x)^b(x+1-x)\ dx\\ =\int_0^1 x^a(1-x)^b\ dx\]

OpenStudy (blockcolder):

Part 3 is straightforward integration by parts: \[\int_0^1 x^{a+1} b(1-x)^{b-1}\ dx=\left .-x^{a+1}(1-x)^b\right |_0^1+\int_0^1 (a+1)x^a(1-x)^b\ dx\\ b\int_0^1 x^{a+1} (1-x)^{b-1}\ dx=(a+1)\int_0^1 x^a(1-x)^b\ dx\]

OpenStudy (anonymous):

correct!

OpenStudy (unklerhaukus):

Reminds me of the Beta function

OpenStudy (anonymous):

i hadn't come across the beta function before, very interesting! thanks

OpenStudy (anonymous):

reading this now http://en.wikipedia.org/wiki/Beta_function

OpenStudy (blockcolder):

I can't decrease a and b at the same time. :(

OpenStudy (anonymous):

do you need to? if we can decrease one to zero or one, the integral is easy